OpenStudy (anonymous):

( x ) | x in R ( 1 ) is it subspace in R^2

OpenStudy (anonymous):

let's do scalar multiple c* (x) = cx (1) c test for linearity no because we need and 1 and we get a c

OpenStudy (anonymous):

@zarkon, can you confirm

OpenStudy (zarkon):

is that the vactor \[\left[\begin{matrix}x \\ 1\end{matrix}\right]\]

OpenStudy (zarkon):

vector

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

subspace contains the zero vector

OpenStudy (zarkon):

your first post also works

OpenStudy (anonymous):

how does zero vector explanation work

OpenStudy (zarkon):

the zero vector is \[\left[\begin{matrix}0\\ 0\end{matrix}\right]\] \[\left[\begin{matrix}x \\ 1\end{matrix}\right]\] is not of that form

OpenStudy (anonymous):

oh , since 1 can't ever be 0

OpenStudy (zarkon):

you could also do this \[\left[\begin{matrix}x \\ 1\end{matrix}\right]+\left[\begin{matrix}y \\ 1\end{matrix}\right]=\left[\begin{matrix}x+y \\ 2\end{matrix}\right]\] which is also not of the original form

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

nice, thanks